Exercise 10-25 Sandhill Co. issued $477,000, 6%, 30-year bonds on January 1, 2022, at 105. Interest is payable annually on JaClick if you would like to Show Work for this question: Open Show Work
Answer #1
Answer
Face Value =$ 477000
Premium =23850
Amortisation of premium =23850/30year=795 per year
Interest on bonds =477000*6% =$ 28620
Interest payable =28620-795= $27825
Date Particulars debit credit
01-Jan-22 Cash $       5,00,850
Bond payable $       4,77,000
Prerimum on bond payable $           23,850
31-Dec-22 Interest Exp $           27,825
Prerimum on bond payable $                 795
Interest Payable $           28,620
01-Jan-23 Interest Payable $           28,620
Cash $           28,620
01-Jan-43 Bond payable $       4,77,000
Cash $       4,77,000